2. How many moles of Na2O (sodium oxide) are produced if 1.5 mol of Fe(iron) is

produced in the following reaction:

6Na + Fe2O3 - 3Na,O + 2Fe

3. If 4.3 moles of CaC2 (calcium carbonide) are used up in the reaction below, how

many grams of HO(water) are needed?

CaC2 + H20 - CaH2 +Ca(OH)?


1
Expert's answer
2022-06-15T11:02:04-0400

Solution (2):

Balanced chemical equation:

6Na + Fe2O3 → 3Na2O + 2Fe

According to stoichiometry:

When 2 mol of Fe are produced, 3 mol of Na2O are produced.

When 1.5 mol of Fe are produced, X mol of Na2O are produced.

Therefore,

Moles of Na2O = X = (1.5 mol Fe) × (3 mol Na2O / 2 mol Fe) = 2.25 mol Na2O

Answer (2): 2.25 moles of Na2O (sodium oxide) are produced.


Solution (3):

Balanced chemical equation:

CaC2 + 2H2O → C2H2 + Ca(OH)2

According to stoichiometry:

1 mol of CaC2 reacts with 2 mol of H2O

Thus, 4.3 mol of CaC2 react with:

(4.3 mol CaC2) × (2 mol H2O / 1 mol CaC2) = 8.6 mol H2O


The molar mass of water (H2O) is 18.0 g/mol

Therefore,

Mass of H2O = (8.6 mol H2O) × (18.0 g H2O / 1 mol H2O) = 154.8 g H2O

Answer (3): 154.8 grams of H2O (water) are needed.